2.5 F of elec is passes thru a soln of CuSo4.the no of gm equivalents of Cu deposited on the cathode will be.?

Grade:12

4 Answers

Naveen Kumar
9 years ago
The deposition of Cu can be given as:
Cu2+............+ 2e-................>Cu(s)
the number of electron involved =2 for making one mole of Cu
As we know that one mole of electron having charge=1F
so by passing, 1 mole of e- or 1F the moles of Cu deposited = 1/2
so for 2.5 mole, moles of Cu deposited=2.5/2 mole=2.5/2 * (Molecular weight of Cu) g
Shriya Mehrotra
9 years ago
sir but the ans is 2.5
Naveen Kumar
9 years ago
The number of Gm-equivalents=Mass of Cu/Equivalent weight
Equivalent equivalent weight=Atomic weight /n-factor
n-factor=number of electrons per one atom of Cu
As we know the mass deposzsited in gram and n-factor=2 and we know the atomic weight so plz use the resultsand proceed further.
britto george
9 years ago
this is a simple question and you can get good things from solving this so just try